Find the maximum displacement response spectrum (normalized to the static displacement) of a damped linear single degree of freedom oscillator with mass=m, stiffness=k, and damping constant=2ζmw0. I believe the correct way to the solution is to use mx"+cx'+kx=0, after dividing both sides by m, the equation becomes mx"+2ζw0x'+wn2x=0, but now I am a little at a loss of what to do next. If the w0 and wn were the same values, then the answer would be easy enough to find. As for the maximum displacement response spectrum, I think that the Duhamel's integral will be used since I already did a problem with it for an undamped case, but any help would be appreciated.
Answer #1

,mx t cx t kx t mx t ++= g    (4.1) Substitute ω0= / k m and 0 2 c m ξ = ω and 2 0 1 ω =ω −ξ d The equation (4.1) can be re-written as 2 0 0 ( ) 2 ( ) ( ) - ( ) g   xt xt xt x t + ξω + ω =  (4.2) Using Duhamel’s integral, the solution of SDOF system initially at rest is given by (Agrawal and Shrikhande, 2006) - ( - ) 0 0 ( ) - ( ) s ( - ) t t g d d e x t x in t d ξω τ = τ ω ττ ω ∫  (4.3) The maximum displacement of the SDOF system having parameters of ξ and ω0 and subjected to specified earthquake motion, ( ) g x t is expressed by - ( - ) 0 max 0 max ( ) ( ) s ( - ) t t g d d e x t x in t d ξω τ = τ ω ττ ω ∫  (4.4) The relative displacement spectrum is defined as, d 0 max S ( , )= ( ) ξ ω x t (4.5) where 0 (, ) d S ξ ω is the relative displacement spectra of the earthquake ground motion for the parameters of ξ and ω0. Similarly, the relative velocity spectrum, Sv and absolute acceleration response spectrum, Sa are expressed as, v 0 max S ( , )= ( ) ξ ω x t  (4.6) a 0 max max S ( , )= ( ) ( ) ( ) a g ξω = +    x t xt x t (4.7) The pseudo velocity response spectrum, Spv for the system is defined as S ( , ) = S ( , ) pv 0 0 d 0 ξω ω ξω (4.8) Similarly, the pseudo acceleration response, Spa is obtained by multiplying the Sd to ω0 2 , thus 2 S ( , ) = S ( , ) pa 0 0 d 0 ξω ω ξω (4.9) Consider a case where 2 0 0 . . ( ) ( ) - ( ) g ξ= +ω = ie x t x t x t   max | ( ) ( )| S xt x t a g = +   2 0 max =−ω | ( )| x t 2 0 max = ω | | x 2 = ω0Sd = Spa (4.10) The above equation implies that for an undamped system, Sa = Spa. The quantity Spv is used to calculate the maximum strain energy stored in the structure expressed as 2 2 2 2 max max 0 11 1 22 2 E kx m S mS = = ω= d pv (4.11) The quantity Spa is related to the maximum value of base shear as 2 max max 0 pa